The manufacturer of a brand of juice drink claims that the mean content of its product is 210 ml with a standard deviation of 6.7 ml.


What is the probability that a randomly selected sample of the product contains more than or equal to 213 ml?


What is the probability that a sample contains less than 200ml?



1
Expert's answer
2021-12-20T07:12:20-0500

Given:

Mean "(\mu)=" 210, Standard deviation "(\sigma)=" 6.7

Let X be the random variable.

Now,

The probability that a randomly selected sample of the product contains more than or equal to 213 ml is

"P(X>213)=P(\frac{X-\mu}{\sigma}>\frac{213-210}{6.7})=P(Z>0.44)=0.32997"

Hence, The probability that a randomly selected sample of the product contains more than or equal to 213 ml is 0.32997.


Now, the probability that a sample contains less than 200ml is

"P(X<200)=P(\frac{X-\mu}{\sigma}<\frac{200-210}{6.7})=P(Z<-1.49)=0.068112"

Hence, the probability that a sample contains less than 200ml is 0.068112.